OpenStudy (anonymous):

9c-6d+4c-3d combine like terms and simplify. will give medal

OpenStudy (anonymous):

I added the terms together and simplified to 13c+9d

TheSmartOne (thesmartone):

13c part is correct but what is -6d - 3d = ?

OpenStudy (anonymous):

-3

TheSmartOne (thesmartone):

umm no -6 -3 = -1(6+3) = ?

OpenStudy (anonymous):

why would we subtract those ?

OpenStudy (anonymous):

I dont understand

TheSmartOne (thesmartone):

-6 -3 factor out -1 so you can see the final answer -6 = -1 x 6 -3 = -1 x 3 -6 - 3 = -1(6) + (-1)(3) = -1(6+3) = ?

jimthompson5910 (jim_thompson5910):

start at -6 

jimthompson5910 (jim_thompson5910):

then go 3 units to the left (we go 3 units left because we're subtracting 3)  so that's why `-6 - 3 = -9`

OpenStudy (anonymous):

I got -9d

TheSmartOne (thesmartone):

correct :)

OpenStudy (anonymous):

would the final equation be 13c+-9d?

TheSmartOne (thesmartone):

so your answer would be 13c - 9d not 13c + 9d

OpenStudy (anonymous):

Thank you

TheSmartOne (thesmartone):

anytime :)